OpenStudy (anonymous):

Find vertical asymptotes and prove with limits of f(x) = ((x^2+x)/(x^2-1))

OpenStudy (anonymous):

\[f(x) = ((x^2+x)/(x^2-1)) \]

OpenStudy (amistre64):

how do we define a vertical asymptote

OpenStudy (anonymous):

as x approaches a number, y value gets closer to infinity or negative infinity

OpenStudy (anonymous):

but why is it that only x=1 is a vertical asymptote and not x = -1

OpenStudy (loser66):

because when x= -1 , f(x) = 0/0, and we must use l'hopital rule to find lim of f(x) which gives us lim =1

OpenStudy (anonymous):

how do i use that rule? can you please explain to me

OpenStudy (loser66):

when you have f(x) of the form of 0/0 or \(\dfrac{\infty}{\infty}\), then you can apply l'hopital rule to find the lim, just take derivative both numerator and denominator , and then take lim of it. If it is of those form, again, take derivative until you can take limit.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Do you actually mean "prove" with the \(\epsilon-\delta\) definition, or do you just need to compute?